# RealCQA-V2 : Visual Premise Proving thanks: WIP: Data will be uploaded soon

### v1

@@ 3.3 VPP Evaluation
The $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency unlike traditional metrics that allow partial correctness. It requires that every premise in a reasoning sequence be validated as correct for the entire sequence to be considered accurate, reflecting the principles of formal logic where any single error can invalidate an entire proof. This follows from the nature of deductive reasoning, where conclusions are only as valid as their weakest premise. The Depth of Correct Premises ($DCP$) metric complements $Acc_{VPP}$ by providing insight into the extent of errors within incorrect sequences.
@@ 3.3.1 Accuracy of Proven Premises
This metric assigns a maximum value of 1 for sequences where all premises are true and 0 if any premise is incorrect. The multiplicative approach reflects the logical "AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true. If a model predicts a final conclusion as True, all intermediate premises in its reasoning sequence must also be True; otherwise, an incorrect intermediate premise indicates that the model relied on an incorrect bias to reach the correct conclusion.
@@ 3.3.2 Depth of Correct Premises
${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
Here, $S_{correct}$ represents the number of sequences where all premises are correct. The $DCP$ metric first identifies incorrect sequences and normalizes the sum of correct premises by the total number of premises in each chain. This provides a measure of how far a model correctly reasons within sequences that are not fully accurate, highlighting areas for potential improvement in model training and reasoning formulation.
While metrics like NDCG and MRR offer measures for structured output, they do not assess the depth of correctness within reasoning sequences. The $DCP$ metric captures how much of the reasoning chain is correct even when the overall sequence is not, providing a nuanced understanding of the model’s capacity to handle extended reasoning tasks.
@@ 4 Experimentation
We conduct all our experiments using a 12 layer encoder-decoder transformer model with a resolution preserving image tokenizer as introduced in Pix2Struct. We employ the variant pretrained with math and chart question answering following Matcha. All our experiments are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2. The model uses a fixed context length of 2048 tokens. And inference with a single text-image pair costs about $\sim$5G memory. ChartPrem model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task. 1 full epoch with chart qa (1.5mil) + challenge json (30k) + structure premises(52k) takes around 2 weeks in our setting. We train for 150k iterations(see supplementary) and use all types of the premises for evaluation (158k) of our proposed task, using a subset of the extensively generated data premises(60k). We hope by publicly releasing the full dataset, models and evaluation code, it will stimulate the community for further exploration especially using the extensive premises for instruction fine-tuning LVLM’s and lead to improvement of models to parse large scale scientific publication repositories.
4.1 Dense Chart Parsing
We compare our model’s performance on dense chart parsing tasks by using structured premises for structure prediction, such as querying ’What is the type of chart?’ or ’What is the title of the dependent axis?’ instead of conventional image classification. We compare with the results of the corresponding tasks from the previous challenge[11], and in the column ‘Direct Prediction’ we report results as reported in the challenge report by task specific vision models. The next column is the Matcha-base model in a zero shot setting, and the next column ‘Matcha-FT’ is the same model when trained only on the RealCQA QA pairs. The next column, ‘ChartPrem’ is our proposed model trained further on the SP’s we created. We provide exhaustive list of our queries in the appendix. These queries are complementary to the underlying binary SP’s that we created for the VPP task and have text answers. Thus, training on SP’s improves performance on previous dense chart parsing tasks as compared to direct pixel level predictions. While zero-shot Matcha is only able to generalize sufficiently to three chart properties categorical labels, logarithmic axis and presence of legend, on fine-tuning the performance improves but the vision based task specific models still outperform. Only on further training with the SP’s we see considerable improvement. The worst performance is for dependent axis title which can include multiple complex math symbols used in scientific charts which the model might not have seen as much due to the text heavy pre-training datasets. The second worst performance is for categorical x-tick labels and this is due to the complex grouped and stacked bar charts which might again have math-symbol intensive labels and at times are quite cluttered having 40-50 tilted labels on a single chart.
Chart Type 94.63 35.21 80.96 98.72 F1 - Precision/recall
Dependent(Y) Axis Title 75.62 21.85 47.82 77.80 Text Matching
Y-Min Value 73.42 34.61 70.1 97.59 Absolute Value
Y-Max Value 62.22 24.74 54.63 96.5 Absolute Value
Independent(X) Axis Title 85.62 42.3 79.34 92.38 Text Matching
X-Min Value 73.42 24.61 62.77 96.94 Absolute Value
X-Max Value 62.22 34.74 57.02 95.31 Absolute Value
Categorical X-Tick Labels 68.83 33.78 58.76 79.74 1:1 Accuracy
Is Categorical or Not - 82.97 91.48 100 Binary
Is Logarithmic or Not - 84.65 93.24 100 Binary
Is Legend Present or Not - 77.32 90.66 100 Binary
Number of Data Series - 38.43 56.57 82.61 Absolute Value
Legend Name 82.93 47.92 73.86 91.27 1:1 Accuracy
@@ 4.2 NLP QA
"We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views. Figure 4 shows the delta performance of each model with comparison to zero-shot matcha-base. These results are presented in Table 3, where the first column describes the category of QA, the rest are performances by different models. These are all reported as accuracy percent over total QA pairs in that category except ranked lists which are evaluated using nDCG@10.
@@
The next categorisation is by question type. It is interesting to note the zero-shot Matcha model has better performance than CRCT on structural and reasoning questions owing to their larger scale pre-training dataset, but lags in retrieval due to the visual domain shift of scientific charts. Both zero shot models Unichart and Matcha perform better on reasoning type questions than structural and retrieval based questions. This suggests these models are capable of leveraging other inductive biases present in the data and perform well without actually performing ‘reasoning’ as depicted by only looking at this metric.
@@
4.3 Visual Premise Proving
We evaluate the ChartPrem model on our VPP metrics.
4.3.1 FOL Solver
4.3.2 Chain of Reasoning Evaluation

### v2

@@ 3.3 VPP Evaluation
We propose two metrics. The Accuracy of Proven Premises $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency. The Depth of Correct Premises ($DCP$) metric complements $Acc_{VPP}$ by providing insight into the extent of errors within incorrect sequences.
@@ 3.3.1 Accuracy of Proven Premises
This metric assigns a maximum value of 1 for sequences where all premises are true and 0 if any premise is incorrect. The multiplicative approach reflects the logical “AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true. If a model predicts a final conclusion as True, all intermediate premises in its reasoning sequence must also be True; otherwise, an incorrect intermediate premise indicates that the model relied on an incorrect bias to reach the correct conclusion.
@@ 3.3.2 Depth of Correct Premises
${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),$
where $\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
Here, $S_{correct}$ represents the number of sequences where all premises are correct. The $DCP$ metric first identifies incorrect sequences and normalizes the sum of correct premises by the total number of premises in each chain. This provides a measure of how far a model correctly reasons within sequences that are not fully accurate thus quantifying how much of the reasoning chain is correct when the overall sequence is not. This provides explanation highlighting areas for potential improvement in model training and/or reasoning formulation.
@@ 4 Experimentation
One set of our experiments uses a small sota chart-question answering model, Matcha (282Mil params)[37]. For baseline we fine-tune the Matcha-base with RealCQA [6] question answers(80%) and the raw ground-truth annotations(20%) having all chart attributes from the original challenge [42]. To showcase the usefulness of our proposed premises for NLP-QA for charts we then fine-tune this model, ChartPrem-S(mall) with the premise conclusions and report results on the base chart question answering task as setup in the original Real-CQA paper. We also demonstrate generalizability by reporting results as zero shot for both these models on ChartAriv dataset. In supplementary we also showcase usefulness of chart premises towards the dense chart attribute prediction task. These are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2. The model uses a fixed context length of 2048 tokens. And inference with a single text-image pair costs about $\sim$5G memory. ChartPrem-S model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task. 1 full epoch with chart qa (1.5mil) + challenge json (30k) + structure premises(52k) takes around 2 weeks in our setting. We train for 150k iterations(see supplementary) and use all types of the premises for evaluation (158k) of our proposed task, using a subset of the extensively generated data premises(60k).
Our next set of experiments are using an open source LVLM the ChatIntern2-8b (8.1b params). We repeat similar setting of benchmarking the publicly available weights, finetuned only on QA from RealCQA and raw annotations and further the ChartPrem-L(arge) model finetuned with our proposed premise conclusion pairs. The ChartPrem-L model is finetune in lora setting on 2xA100(80g) GPU’s. We use the authors default setting and also train the vision encoder. We report results for VPP and QA.
We will first discuss the VPP task analysis and then come back to the vanilla chart question answering results.
Figure 7: VPP Analysis
4.1 Visual Premise Proving
We evaluate the ChartPrem model on our VPP metrics.
4.1.1 FOL Solver
4.1.2 Chain of Reasoning Evaluation
@@ 4.2 NLP QA
We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views. These results are presented in Table 1, where the first column describes the category of QA, the rest are performances by different models. These are all reported as accuracy percent over total QA pairs in that category except ranked lists which are evaluated using nDCG@10.
@@
Models Accuracy FOLS $ACC_{VPP}$ DCP # Model Params
ChartPrem 0.46 0.2 0.395
Llava 0.88 0.34 0.73
Gemini 0.56 0.21 0.43
GPT-4o 0.27 0.14 0.38
@@
The next categorization is by question type. It is interesting to note the zero-shot Matcha model has better performance than CRCT on structural and reasoning questions owing to their larger scale pre-training dataset, but lags in retrieval due to the visual domain shift of scientific charts. Both zero shot models Unichart and Matcha perform better on reasoning type questions than structural and retrieval based questions. This suggests these models are capable of leveraging other inductive biases present in the data and perform well without actually performing ‘reasoning’ as depicted by only looking at this metric.